How many grams of ice at -24.9 degrees Celsius can be completely converted to liquid at 5.1 degrees Celsius if the available heat for this process is 4.22x10^3 kJ?

For ice, use a specific heat of 2.01J/(gxC) and enthalpy of fusion of 6.01 kJ/mol.
Answer #1

Answert Given that, Initial temperatude --24.9°C. Final temperatude - 5,1°C. ? = 4022x103KJ. Specific heat -2.01.5/99 off: 60.6.22X103 KJ = 50.04&J + 3 3430+21.3x 4122X100 RT = $450.04 +334+ 21.3) 7 4:22 x103 K9 = 405.34 x J * = 4.22 X10° KJ . 405.34